OpenStudy (anonymous):

Which ordered pairs make the equation true 3x + 2y = –7 (–3, 1) (–2, –1) (1, –4) (3, –8)

OpenStudy (anonymous):

To solve this, you would have to take the answers, let's use (-3, 1) and plug them into the equation: \[x = -3, y = 1\] \[3(-3) + 2(1) = -7\] Now, which of them do you think is the answer?

OpenStudy (anonymous):

i have never done this hahah i have no clue! what are you doing with x+-3,y=1 sorry im not the brightest

OpenStudy (anonymous):

how do i solve the bottem one? 3(-3) + 2(1) = -7

OpenStudy (anonymous):

@Nibby

OpenStudy (anonymous):



OpenStudy (anonymous):

can you explain how to solve this? you dont have to gve me the answer just explain it?

OpenStudy (anonymous):

To solve the bottom one, times 3 by -3, which gives you -9. Then, times the 2 by 1, which gives you 2. Now you add them together, -9 + 2 = ?

OpenStudy (anonymous):

-7! omg your a life saver!!

OpenStudy (anonymous):

Sure. The answers are given in X and Y, So let's look at the first answer. The first answer is (-3, 1). The number on the left is the X and the number on the right is Y. So, X = -3, and Y = 1. Since we know this, and the equation is asking: 3x + 2y = -7, then we can plug in the X and the Y into the equation. Since we know X is -3, and Y is 1, we can then change the equation to: 3(-3) + 2(1) = -7 Make sense?

OpenStudy (anonymous):

yep do i always multiply

OpenStudy (anonymous):

If the equation is 3 + x, then you would put X in. So, if our X is still -3 (It won't always be the same!) the equation becomes: 3 + (-3)

OpenStudy (anonymous):

So, no, you will not always multiply.

OpenStudy (anonymous):

so x equales multiply ?

OpenStudy (anonymous):

No, no. X is the substitute of a number we don't know. So, in the equation: \[3x + 1 = 0\] We don't know what X is yet. We have to find it. X does not mean multiply, it's just a number we don't know yet.

OpenStudy (anonymous):

ahhh ok

OpenStudy (anonymous):

can u help me set up the next one. just the basic first 2 steps

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

just for practice the second option on this questopn(–2, –1)

OpenStudy (anonymous):

ugh (-2,-1)

OpenStudy (anonymous):

Since our equation is: \[3x + 2= -7\] We have to take the X and Y from (-2, -1) and plug them in. Since the first number in the pair is X, we now know that X is -2, and Y is -1, since Y is the second number in the pair. Since we know this, we can plug them straight into the equation, so our equation becomes: \[3(-2) + 2(-1) = -7\] In order to solve this, we times 3 by -2, because 3(-2) is basically 3 * (-2). This becomes -6. Next, we have to times 2 by -1, because 2(-1) is 2 * (-1). This becomes -2. Then we plug them in, so our equation becomes: \[-6 + - 2 = -7\] We know this can't be the answer because -6 + -2 = -8, not -7.

OpenStudy (anonymous):

@Nibby

OpenStudy (anonymous):

x + 3y < 9 is it supposed to equal 9 or less than 9

OpenStudy (anonymous):

The above statement means that 9 is greater than x + 3y. Because < Means the number on the right is greater, > means the number on the right less, and = means they are equal. Example: 1 < 2 2 > 1 2 = 2

OpenStudy (anonymous):

yes. i know that but what is this equation asking for? would my naswer equal 9 or ?

OpenStudy (anonymous):

@Nibby

OpenStudy (anonymous):

The x + 3y is less than 9.